Consider a small open economy (e.g. the Netherlands) producing two goods, clothing and food. The clothing industry uses capital (K) and labor (LC) as inputs, while the food industry uses land (La) and labor (LF ) as factors of production. The production technologies for the two industries are given by QC = K ¼ LC 3/4 ; QF = La1/2L F 1/2 . Also, the country is endowed with 216 units of capital, 360 units of labor, and 9 units of land. Finally, the utility function of the representative consumer is given by U(CC, CF ) = CF 1/2Cc 1/2 .

1.Calculate the autarky production (= consumption) for the two industries. 2. Derive the autarky price ratio, ( pC/pF ) and the real value of production (i.e. real GDP) in autarky. 3. Determine the autarky utility level. 4. Determine LC and LF in autarky. 5. Calculate the realwage rate ( w/pF ), the real interest rate ( r/pF), and the real return on land ( i /pF) in autarky. 6. Calculate real total factor income in autarky. Is it equal to the real value of production that you calculated previously? 7. Now, assume that the country is allowed to trade with the rest of the world at the free trade price ratio, ( pC/pF ) , equal to 2/3 . Determine LC and LF in freetrade. 8. Calculate the real wage rate, the real interest rate, and the real return on land in free trade. 9. Who are the winners and losers of the move from autarky to trade? 1 10. Calculate real total factor in come in freetrade. 11. Calculate free trade consumption of food and clothing. 12. Determine the level of utility with free trade. Are the Dutch consumers betteroff?
Answer #1

1. Autarky production is the self-sufficient production of an economy

units given:

k = 216

LC = 360

La = 9

LF = 360

Production techniques for clothing industry

= 216 * (1/4 * 360) * 3/4 = 14580

Production techniques for food industry

= 9 * (1/2 * 360) * 1/2 = 810

2. Autarky price ratio (food over clothing) is the opportunity cost of producing food